OpenStudy (anonymous):

Would you make more interest by investing five thousand dollars in a 6-month certificate of deposit (CD) at six percent interest or by investing three thousand dollars in a twelve-month CD at four percent interest? Show your work.

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (arabpride):

tbh i searched that question nd there where PLENTYY of links for the answer so ur answer is gonna be on google *^.^*

OpenStudy (whpalmer4):

You don't specify the compounding, so I'll assume simple interest. \[I = Prt\] where \(I\) is the interest earned or paid, \(P\) is the principal, \(r\) is the interest rate expressed as a decimal, and \(t\) is the time in years \[I_{5k} = 5000*0.06*\frac{1}{2}=\]\[I_{3k} = 3000*0.04*1=\]